OpenStudy (anonymous):

multiply or divide. 4x+12/x^2-2x times x/6x+18 help!!!

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

(4x+12)/(x^2-2x) = (4(x+3))/(x(x-2)), x/(6x+18) = x/(6(x+3)), multiply them (4(x+3))/(x(x-2)) * (x/(6x+18) ... cancel out what u can, not sure if this helps

OpenStudy (anonymous):